# Renormalizable Spectral-Shell Dynamics as the Origin of Neural Scaling Laws

### v4

@@ 1 Introduction
Modern deep neural networks are trained by strongly nonlinear and high-dimensional optimization procedures. Nevertheless, across architectures, datasets, and training recipes, their performance exhibits remarkably simple and robust regularities. Most prominently, empirical studies have shown that training and test losses obey approximate power laws in model size, dataset size, or total compute over wide dynamic ranges (hestness2017deep; kaplan2020scaling; henighan2020scaling; hoffmann2022training; hernandez2021scaling; wei2022emergent). Related law-like behaviors have been observed in precision scaling, pruning, sparsification, and model densification (kumar2024scaling; sorscher2022beyond; rosenfeld2021predictability; blalock2020state; Xiao2024DensingLaw). These phenomena strongly suggest the existence of an underlying macroscopic structure governing how error is redistributed and dissipated during training.
@@
A classical theoretical explanation is provided by the neural tangent kernel (NTK) framework (jacot2018ntk; lee2019wide; bietti2021inductive), which models training as a linear dynamics in function space with a fixed kernel. While this perspective explains certain spectral biases and convergence rates, it relies on a frozen feature map and breaks down once representation learning becomes significant. Recent work has therefore emphasized feature-learning dynamics, using mean-field, tensor-program, or dynamical systems approaches (yang2021tensor; canatar2022kernel; bordelon2024dynamical; bordelon2023feature; wang2023spectralevolution). These models successfully capture richer behavior beyond the lazy regime, but typically treat NTK-like and feature-learning regimes via distinct approximations and do not provide a single operator-level equation that interpolates continuously between them.
@@ 3.1 Exact Mode ODEs in a Drifting Eigenbasis
The second term encodes the rotation of the eigenbasis. From Kato’s perturbation theory (kato2012short; Zwiebach_Adiabatic) for differential self-adjoint operators, the evolution of the eigenfunctions satisfies
@@ 4.3 Definition: renormalizable shell dynamics via cumulative energy flux
The notion of renormalizability adopted here follows the standard usage in statistical physics and field theory. Rather than requiring microscopic locality or exact continuum limits, external interaction effects on a resolution shell are integrated out and absorbed into effective inter-shell fluxes, while subleading corrections become irrelevant at coarse scales. This philosophy underlies Wilsonian renormalization group theory (wilson1983renormalization; kadanoff1966scaling), shell models of turbulence and energy cascades (kolmogorov1995turbulence), and effective hydrodynamic descriptions of nonequilibrium systems (forster1977large; spohn2012large). Our definition formalizes this principle at the level of spectral-shell energy dynamics.
@@ 4.4 PDE-approximability across shells (energy continuum)
This is a local continuum approximation over that range; we do not require a global continuous-spectrum limit. As in nonequilibrium statistical physics forster1977large; spohn2012large, we do not assume the existence of an exact continuum limit; rather, renormalizability refers to the existence of a closed effective description after coarse-graining. We stress again that our continuum equations do not assume the existence of a genuine continuous-spectrum limit. Under this approximation we acquire the expression of the total loss function:
@@ 5 Scale-Free GRSD Under Power-Law–Compatible Transport
At the discrete shell level, renormalizability asserts that the cumulative effect of inter-shell interactions can be summarized by a controlled boundary flux. In the continuum approximation, this structure manifests as a transport term that can be absorbed into a time-dependent reparameterization of the spectral coordinate. The characteristic flow (5) implements precisely this absorption.
@@ 5.2 Energy density along characteristics
which is aligned with the spectral evolution conjecture proposed in zhang2025generalized:
@@
where $g(\lambda,t)\propto\lambda^{a(\beta)}t^{b(\beta)}$ and thus $w_{\lambda}^{2}$ correspond to the starting error on mode with eigenvalue of $\lambda$. This conjecture is helpful for understanding the scaling law for both loss evolution and model compression. It further predicted the existence of learning frontier and defines model density, which is observed to evolve according to Xiao2024DensingLaw.
@@
$\mathcal{L}(t)=\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda.$
Substituting (13) and performing a standard moving-front estimate yields power-law decay
$\mathcal{L}(t)\sim t^{-(1+c_{0})(1-\alpha)},$
for initial spectra $\varepsilon_{0}(u)\sim u^{-\alpha}$.
@@ 6 Double Descent from In-Domain Finite-Sample Spectral Mismatch
For continuous data distributions or infinite-rank kernel operators, the population spectrum exhibits vanishing eigenvalue gaps in the spectral tail. Classical results on empirical covariance and integral operators show that, in this regime, the associated empirical eigenspaces exhibit $O(1)$ fluctuations under finite sampling (koltchinskii2017normal; rosasco2010learning). Consequently, even under in-domain sampling, $\Delta$ is generically supported on high-resolution (spectral-tail) components and is not co-diagonalizable with the training-induced operator $M(t)$.
@@ 7 Related Work
Empirical studies have established that neural networks display remarkably regular power-law relationships between compute, model size, dataset size, and achievable loss (kaplan2020scaling; henighan2020scaling; hestness2017deep; hoffmann2022training; hernandez2021scaling). Recent work has investigated both the emergence of such scaling behavior and its theoretical underpinnings, including dynamical models of loss evolution (bordelon2024dynamical) and the role of data pruning (sorscher2022beyond). Precision scaling laws have also been explored in the context of architectural and quantization constraints (kumar2024scaling).
@@
A major line of work studies how neural networks acquire hierarchical representations and exhibit spectral preference (rahaman2019spectral; bordelon2023feature). Kernel-based analyses of learning dynamics (bietti2021inductive; canatar2022kernel) and infinite-width approximations (jacot2018neural; lee2019wide; yang2021tensor) have contributed significantly to understanding the transition between lazy training and representation learning. Recent studies have also revealed consistency of learned features across widths (vyas2023feature) and the spectral evolution of networks (wang2023spectralevolution).
@@
The geometry of optimization landscapes, the effect of batch size, and training stability have been examined extensively (keskar2017large; ghorbani2019investigation). Work on pruning and compression (rosenfeld2021predictability; han2015deep; blalock2020state; lecun1990optimal; molchanov2017variational; lee2019snip; wang2020picking; han2016deep; nagel2021up; frantar2022gptq; dettmers2022llm) has illuminated how spectral structure interacts with parameter sparsity and low-precision computation. Theoretical connections between spectral evolution, implicit bias, and neural dynamics continue to be an active area of research (domine2024lazy; zhang2025generalized).
@@
Our spectral formulation draws on classical operator perturbation theory (kato2012short) and its analogues in quantum adiabatic evolution (Zwiebach_Adiabatic). These tools formalize how eigenfunctions and eigenvalues evolve under smooth or weakly coupled updates, providing a principled foundation for the drift–dissipation dynamics developed in this work.

### v5

@@ 1 Introduction
Modern deep neural networks are trained by strongly nonlinear and high-dimensional optimization procedures. Nevertheless, across architectures, datasets, and training recipes, their performance exhibits remarkably simple and robust regularities. Most prominently, empirical studies have shown that training and test losses obey approximate power laws in model size, dataset size, or total compute over wide dynamic ranges (Hestness et al., 2017; Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022; Hernandez et al., 2021; Wei et al., 2022). Related law-like behaviors have been observed in precision scaling, pruning, sparsification, and model densification (Kumar et al., 2024; Sorscher et al., 2022; Rosenfeld et al., 2021; Blalock et al., 2020; Xiao et al., 2025). These phenomena strongly suggest the existence of an underlying macroscopic structure governing how error is redistributed and dissipated during training.
@@
A classical theoretical explanation is provided by the neural tangent kernel (NTK) framework (Jacot et al., 2018b; Lee et al., 2019a; Bietti and Bach, 2021), which models training as a linear dynamics in function space with a fixed kernel. While this perspective explains certain spectral biases and convergence rates, it relies on a frozen feature map and breaks down once representation learning becomes significant. Recent work has therefore emphasized feature-learning dynamics, using mean-field, tensor-program, or dynamical systems approaches (Yang, 2021; Canatar and Pehlevan, 2022; Bordelon et al., 2024; Bordelon and Pehlevan, 2024; Wang et al., 2023). These models successfully capture richer behavior beyond the lazy regime, but typically treat NTK-like and feature-learning regimes via distinct approximations and do not provide a single operator-level equation that interpolates continuously between them.
@@ 3.1 Exact Mode ODEs in a Drifting Eigenbasis
The second term encodes the rotation of the eigenbasis. From Kato’s perturbation theory (Kato, 2012; Zwiebach, 2018) for differential self-adjoint operators, the evolution of the eigenfunctions satisfies
@@ 4.3 Definition: renormalizable shell dynamics via cumulative energy flux
The notion of renormalizability adopted here follows the standard usage in statistical physics and field theory. Rather than requiring microscopic locality or exact continuum limits, external interaction effects on a resolution shell are integrated out and absorbed into effective inter-shell fluxes, while subleading corrections become irrelevant at coarse scales. This philosophy underlies Wilsonian renormalization group theory (Wilson, 1983; Kadanoff, 1966), shell models of turbulence and energy cascades (Kolmogorov, 1995), and effective hydrodynamic descriptions of nonequilibrium systems (Forster et al., 1977; Spohn, 2012). Our definition formalizes this principle at the level of spectral-shell energy dynamics.
@@ 4.4 PDE-approximability across shells (energy continuum)
This is a local continuum approximation over that range; we do not require a global continuous-spectrum limit. As in nonequilibrium statistical physics Forster et al. (1977); Spohn (2012), we do not assume the existence of an exact continuum limit; rather, renormalizability refers to the existence of a closed effective description after coarse-graining. We stress again that our continuum equations do not assume the existence of a genuine continuous-spectrum limit. Under this approximation we acquire the expression of the total loss function:
@@ 5 Scale-Free GRSD Under Power-Law–Compatible Transport
On the origin of the scale-free drift form.
The specific choice $v(\lambda,t)=c_{0}\,\lambda/t$ should not be interpreted as an ad hoc modeling assumption. Under a set of strong but practically natural conditions—such as controlled training trajectories, uniformly bounded gradients, and zero-mean random initialization—the same scale-free drift form can be derived directly from the training dynamics. A concrete derivation under such conditions is given in Zhang (2026). In the present work, we do not rely on these sufficient conditions and instead take Eq. (5) as a macroscopic consequence of scale-free spectral evolution; the aforementioned derivation serves only to demonstrate that the assumed drift is dynamically realizable rather than postulated.
@@
At the discrete shell level, renormalizability asserts that the cumulative effect of inter-shell interactions can be summarized by a controlled boundary flux. In the continuum approximation, this structure manifests as a transport term that can be absorbed into a time-dependent reparameterization of the spectral coordinate. The characteristic flow Eq. (5) implements precisely this absorption.
@@ 5.2 Energy density along characteristics
which is aligned with the spectral evolution conjecture proposed in Zhang (2025):
@@
where $g(\lambda,t)\propto\lambda^{a(\beta)}t^{b(\beta)}$ and thus $w_{\lambda}^{2}$ correspond to the starting error on mode with eigenvalue of $\lambda$. This conjecture is helpful for understanding the scaling law for both loss evolution and model compression. It further predicted the existence of learning frontier and defines model density, which is observed to evolve according to Xiao et al. (2025).
@@
At the level of logarithmic spectral shells, the high-resolution behavior of the quadratic error energy can be expressed in a factorized form
$E_{\alpha}(t)\;\approx\;A(\lambda_{\alpha}\,t^{-c_{0}})\,\exp\!\big(-\kappa(t)\,\lambda_{\alpha}\big),\qquad\lambda_{\alpha}\ \text{beyond the resolution frontier},$ (14)
where $\lambda_{\alpha}$ denotes the representative eigenvalue of shell $\alpha$, $\kappa(t)$ is an effective dissipation scale, and $A(\lambda,t)$ is a coarse-grained amplitude capturing the combined effects of spectral transport and initialization.
Remark (Amplitude structure).
The function $A(\lambda,t)$ is not assumed to take any specific form in general. Particularly, when the initial error spectrum is approximately scale-free over a range of resolutions, the amplitude may inherit a corresponding power-law dependence on $\lambda$:
$\varepsilon(\lambda,t)=\big(\lambda\,t^{-c_{0}}\big)^{p}\,\exp\!\Big(-\frac{2\lambda}{1+c_{0}}\,t\Big)=t^{-c_{0}p}\big(\lambda\big)^{p}\,\exp\!\Big(-\frac{2\lambda}{1+c_{0}}\,t\Big).$ (15)
which leads to a shell-energy profile of the representative form
$E_{\alpha}(t)\;\sim\;C(t)\lambda_{\alpha}^{\,p+1}\,\exp\!\big(-\kappa(t)\,\lambda_{\alpha}\big),\qquad\lambda_{\alpha}\ \text{beyond the frontier},$ (16)
for some exponent $p$. We emphasize that this expression is provided only as an illustrative example of how scale-free initialization can manifest at the shell level; no assumption on the sign or universality of $p$ is made, and our empirical analysis does not rely on this specific functional form.
@@
$\mathcal{L}(t)=\int_{\lambda_{\min}}^{\lambda_{\max}}\varepsilon(\lambda,t)d\lambda.\approx\int_{\lambda_{\min}}^{\infty}\varepsilon(\lambda,t)d\lambda.$ (17)
where the approximation holds for an exponential decay as lambda increases. Starting from
$\varepsilon(\lambda,t)=(\lambda t^{-c_{0}})^{p}\exp\!\Big(-\frac{2\lambda}{1+c_{0}}\,t\Big)=t^{-c_{0}p}\,\lambda^{p}\,e^{-at\lambda},\qquad a:=\frac{2}{1+c_{0}},$ (18)
the total quadratic error is
$L(t)\approx\int_{\lambda_{\min}}^{\infty}\varepsilon(\lambda,t)\,d\lambda=t^{-c_{0}p}\int_{\lambda_{\min}}^{\infty}\lambda^{p}e^{-at\lambda}\,d\lambda.$ (19)
Performing the change of variables $u=at\lambda$, we obtain
$L(t)=t^{-c_{0}p}\,(at)^{-(p+1)}\int_{at\lambda_{\min}}^{\infty}u^{p}e^{-u}\,du=t^{-c_{0}p}\,(at)^{-(p+1)}\Gamma(1+p,at\lambda_{\min}).$ (20)
In our theoretic analysis, the time $t\in(t_{0},T)$ is in the scale-free range. For any $T$ such that $T\lambda_{\min}\ll\frac{1}{a}$, the remaining integral converges to a constant as $t$ varies, yielding the asymptotic scaling
$L(t)\;\propto\;t^{-c_{0}p}\,t^{-(p+1)}=t^{-\big[(1+c_{0})p+1\big]}.$ (21)
@@ 6 Double Descent from In-Domain Finite-Sample Spectral Mismatch
For continuous data distributions or infinite-rank kernel operators, the population spectrum exhibits vanishing eigenvalue gaps in the spectral tail. Classical results on empirical covariance and integral operators show that, in this regime, the associated empirical eigenspaces exhibit $O(1)$ fluctuations under finite sampling (Koltchinskii and Lounici, 2017; Rosasco et al., 2010). Consequently, even under in-domain sampling, $\Delta$ is generically supported on high-resolution (spectral-tail) components and is not co-diagonalizable with the training-induced operator $M(t)$.
@@ 6.4 Double descent from transport-induced amplification in the mismatch window
7 Empirical Spectral-Shell Analysis of LLM Training
This section empirically examines whether the coarse-grained spectral-shell structures assumed by the GRSD framework are observable in realistic large language model (LLM) training checkpoints. Our goal is not to infer microscopic transport mechanisms or fit scaling exponents, but to validate the existence of the shell-level structures required by the theory.
7.1 Experimental setup and error definition
Models and checkpoints.
We analyze publicly released checkpoints from the Pythia model family, including models with 70M, 410M, and 7B parameters. Following the standard Pythia release protocol, checkpoints are provided at regular intervals of approximately 1k optimization steps. All checkpoints analyzed in this section are taken directly from the official releases; no selection based on loss values or convergence criteria is performed.
Error observable.
Let $q_{\theta}(y\mid x)$ denote the model-predicted conditional distribution. Throughout this section, the error signal is defined as the negative log-probability of the ground-truth token,
$e(x)\;:=\;-\log q_{\theta}(y^{*}\mid x),$ (24)
and all spectral quantities are computed from the squared error $e(x)^{2}$.
Justification via quadratic approximation of KL.
For language modeling with one-hot (or near one-hot) supervision, the ground-truth conditional distribution satisfies $p(y^{*}\mid x)\approx 1$. In this setting, the KL divergence reduces to the negative log-likelihood,
$\mathrm{KL}(p\|q_{\theta})=-\log q_{\theta}(y^{*}\mid x)+\mathrm{const}.$ (25)
Moreover, when $q_{\theta}(y^{*}\mid x)\approx 1$, a second-order expansion yields the local quadratic approximation
$\mathrm{KL}(p\|q_{\theta})\;\approx\;\tfrac{1}{2}\,\big(-\log q_{\theta}(y^{*}\mid x)\big)^{2}.$ (26)
Accordingly, the squared log-probability provides a locally equivalent quadratic error measure, consistent with the mean-squared-error framework underlying the spectral-shell analysis. A more detailed derivation is provided in Appendix A.
7.2 Spectral estimation via stochastic Lanczos
Motivation.
The operator $M=J_{\theta}J_{\theta}^{*}$ governing error dynamics acts on a high-dimensional function space, making explicit eigendecomposition infeasible for modern LLMs. To estimate shell-resolved spectral quantities at scale, we employ a stochastic Lanczos method, which probes the spectrum of $M$ using operator–vector products without explicitly forming $M$.
Shell-energy estimation.
The Lanczos procedure produces a low-dimensional tridiagonal approximation whose spectral measure approximates the projection of the error onto the eigenmodes of $M$. The resulting spectral weights are aggregated into logarithmic bins in the spectral coordinate $\lambda$, yielding the log-shell–integrated quadratic error energies
$E_{\alpha}\;=\;\sum_{u:\,\lambda_{u}\in[\lambda_{\alpha},\lambda_{\alpha+1})}g_{u}^{2},$ (27)
which serve as the primary observables throughout this section. All reported curves correspond to $\log E_{\alpha}$ plotted against the shell centers $\log\lambda_{\alpha}$.
Data-induced averaging.
To obtain stable shell-level estimates, all spectral quantities are averaged over multiple independent data batches. Specifically, each estimate aggregates results from approximately 50 batches, each consisting of 128 context windows with batch size 4, corresponding to roughly $2.5\times 10^{5}$ tokens in total. No additional stochasticity is introduced in the Lanczos procedure itself; all averaging arises from the data distribution.
Implementation details and pseudocode for the Lanczos-based shell-energy estimation are provided in Appendix B.
7.3 Shell-energy structure at a fixed checkpoint
(a) Shell-energy profile at a fixed checkpoint.
(b) Piecewise fit with smooth blending.
(c) Minimum shell energy.
Figure 1: Spectral-shell structure of log-probability error for the 7B Pythia model at a representative checkpoint. (a) Minimum energy within each log shell exhibits an abrupt collapse beyond a finite resolution, indicating a sharp dissipation-dominated cutoff rather than an extended power-law tail. (b) A piecewise model—log-linear (power-law) behavior on the low-resolution side and exponential suppression beyond the spectral frontier—provides a good qualitative description of the observed profile. The two regimes are combined via a smooth log-space blending for visualization. (c) The full log-shell energy distribution forms a robust unimodal profile, with a well-defined spectral frontier separating transport-dominated and dissipation-dominated regimes.
Theoretical expectation.
As derived in Eq 16, the shell-integrated quadratic error energy admits a factorized form beyond a finite resolution frontier,
$E_{\alpha}(t)\;\approx\;A(\lambda_{\alpha},t)\,\exp\!\big(-\kappa(t)\,\lambda_{\alpha}\big),\qquad\lambda_{\alpha}\ \text{beyond the frontier},$ (28)
where $A(\lambda,t)$ is a coarse-grained amplitude inherited from transport and initial conditions. In particular, when the initial error spectrum is approximately scale-free over a range of resolutions, the amplitude $A(\lambda,t)$ may exhibit an effective power-law dependence on $\lambda$ over that range.
These observations imply the existence of a well-defined resolution frontier separating two qualitatively distinct regimes. On the low-$\lambda$ side of the frontier, the shell-integrated error energy varies smoothly across logarithmic shells and is well approximated by an affine trend in log–log coordinates, indicating a slowly varying, transport-dominated structure. As $\lambda$ increases, the energy profile reaches a single pronounced maximum, beyond which the behavior changes abruptly.
At fixed training checkpoints, the log-shell energy $E_{\alpha}$ consistently exhibits a unimodal distribution across resolutions: energy accumulates gradually from low $\lambda$, peaks at an intermediate scale, and then decays rapidly at higher resolutions. To probe the high-resolution regime more directly, we examine the minimum error energy within each log shell. Beyond a finite resolution threshold, the minimum shell energy collapses sharply to a numerical floor, producing a clear cutoff on the right side of the spectrum. This abrupt suppression is incompatible with any extended power-law tail and instead indicates a dissipation-dominated regime past the spectral frontier.
7.4 Comparison across checkpoints: spectral-front displacement
(a) Raw log-shell energy profiles.
(b) Smoothed log-shell energy profiles.
Figure 2: Comparison of log-shell error energy distributions across training checkpoints (1k, 2k, 4k, and 8k steps) for a fixed model. (a) Raw shell-energy profiles exhibit a consistent unimodal shape, while the location of the spectral peak and the extent of the high-resolution tail vary systematically across checkpoints. (b) Smoothed profiles suppress shell-level fluctuations and more clearly reveal a leftward displacement of the spectral peak and a contraction of the high-resolution tail as training progresses. Across checkpoints, the low-$\lambda$ side of the spectrum remains approximately parallel in log–log coordinates, indicating coherent reweighting of shell energies rather than independent decay at fixed resolutions.
We next compare log-shell energy profiles across multiple training checkpoints for a fixed model. As shown in Fig. 2, the shell-energy distributions at different checkpoints all retain a pronounced unimodal shape, indicating a stable coarse-grained spectral organization throughout training. However, the location of the peak and the extent of the high-resolution tail shift systematically across checkpoints.
In particular, as training progresses, the position of the spectral peak moves toward smaller $\lambda$, and the high-resolution side of the distribution contracts accordingly. This behavior is more clearly revealed in the smoothed profiles, which suppress shell-level noise while preserving the global shape. The observed shift reflects a gradual displacement of the effective resolution frontier rather than the emergence of qualitatively new spectral features.
At the same time, the low-$\lambda$ side of the spectrum exhibits coherent reweighting across checkpoints. Over a broad range of low resolutions, the shell-energy curves remain approximately parallel in log–log coordinates, while their overall amplitudes increase with training. This indicates that error energy is redistributed across spectral shells in a correlated manner, rather than decaying independently at fixed resolutions.
7.5 Cross-model consistency across scales
(a) Raw log-shell energy profiles.
(b) Smoothed log-shell energy profiles.
Figure 3: Cross-model comparison of log-shell error energy distributions for Pythia models with 70M and 7B parameters at the same checkpoint index. (a) Raw shell-energy profiles show that both models exhibit a similar unimodal structure, with energy rising from low $\lambda$, peaking at an intermediate resolution, and decaying at higher resolutions. (b) Smoothed profiles suppress shell-level fluctuations and highlight the close alignment of the two distributions across a broad range of resolutions. Differences between model scales appear primarily as overall amplitude shifts and modest displacements of the spectral peak, while the low-$\lambda$ side of the spectrum remains approximately parallel in log–log coordinates.
We finally compare shell-energy profiles across models of different scales. Fig. 3 shows the log-shell error energy distributions for Pythia models with 70M and 7B parameters at the same checkpoint index. Despite the substantial difference in model size, the two profiles exhibit strikingly similar coarse-grained structure.
In both cases, the shell-energy distribution forms a pronounced unimodal shape, with energy rising smoothly from low $\lambda$, peaking at an intermediate resolution, and decaying at higher resolutions. The smoothed profiles further highlight this similarity: across a broad range of resolutions, the two curves closely track each other, differing primarily by an overall amplitude shift and a modest displacement of the peak location.
Notably, the low-$\lambda$ side of the spectrum remains approximately parallel in log–log coordinates across model scales, indicating that the relative distribution of error energy across coarse spectral shells is largely preserved. Differences between the 70M and 7B models manifest mainly in the overall scale of energy and the precise position of the spectral peak, rather than in the emergence of qualitatively new spectral features.
This cross-model consistency suggests that the observed shell-level organization reflects a scale-robust property of training dynamics, rather than an artifact of a particular model size.
7.6 Summary of empirical observations
In summary, our experiments establish the following empirical facts: (i) squared log-probability error admits a unimodal shell-energy profile at fixed checkpoints; (ii) a sharp high-resolution frontier exists beyond which shell energy is rapidly suppressed; (iii) the frontier position and low-resolution amplitudes vary coherently across checkpoints; and (iv) these shell-level structures persist across model scales from 70M to 7B parameters.
These observations support the structural assumptions underlying renormalizable spectral-shell dynamics without relying on any specific microscopic transport model.
@@ 8 Related Work
Empirical studies have established that neural networks display remarkably regular power-law relationships between compute, model size, dataset size, and achievable loss (Kaplan et al., 2020; Henighan et al., 2020; Hestness et al., 2017; Hoffmann et al., 2022; Hernandez et al., 2021). Recent work has investigated both the emergence of such scaling behavior and its theoretical underpinnings, including dynamical models of loss evolution (Bordelon et al., 2024) and the role of data pruning (Sorscher et al., 2022). Precision scaling laws have also been explored in the context of architectural and quantization constraints (Kumar et al., 2024).
@@
A major line of work studies how neural networks acquire hierarchical representations and exhibit spectral preference (Rahaman et al., 2019; Bordelon and Pehlevan, 2024). Kernel-based analyses of learning dynamics (Bietti and Bach, 2021; Canatar and Pehlevan, 2022) and infinite-width approximations (Jacot et al., 2018a; Lee et al., 2019a; Yang, 2021) have contributed significantly to understanding the transition between lazy training and representation learning. Recent studies have also revealed consistency of learned features across widths (Vyas et al., 2023) and the spectral evolution of networks (Wang et al., 2023).
@@
The geometry of optimization landscapes, the effect of batch size, and training stability have been examined extensively (Keskar et al., 2017; Ghorbani et al., 2019). Work on pruning and compression (Rosenfeld et al., 2021; Han et al., 2016b; Blalock et al., 2020; LeCun et al., 1990; Molchanov et al., 2017; Lee et al., 2019b; Wang et al., 2020; Han et al., 2016a; Nagel et al., 2021; Frantar et al., 2022; Dettmers et al., 2022) has illuminated how spectral structure interacts with parameter sparsity and low-precision computation. Theoretical connections between spectral evolution, implicit bias, and neural dynamics continue to be an active area of research (Dominé et al., 2024; Zhang, 2025).
@@
Our spectral formulation draws on classical operator perturbation theory (Kato, 2012) and its analogues in quantum adiabatic evolution (Zwiebach, 2018). These tools formalize how eigenfunctions and eigenvalues evolve under smooth or weakly coupled updates, providing a principled foundation for the drift–dissipation dynamics developed in this work.